OpenStudy (anonymous):

is it possible integrate (sinx/x) from 0 to infinity without using engineering math????

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty} \frac{\sin x}{x} dx\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are u familiar with laplace transform?

OpenStudy (anonymous):

yes !!

OpenStudy (anonymous):

\[\int\limits_{0}^{\infty} \frac{\sin x}{x} dx= \int\limits_{0}^{\infty} \int\limits_{0}^{\infty} e^{-xy} dy \sin x \ dx=\int\limits_{0}^{\infty} \int\limits_{0}^{\infty} e^{-xy} \sin x \ dy dx=\int\limits_{0}^{\infty} \int\limits_{0}^{\infty} e^{-xy} \sin x \ dx dy\] intervals are equal so dx dy=dy dx now u have (using laplace transform) \[\int\limits_{0}^{\infty} \int\limits_{0}^{\infty} e^{-xy} \sin x \ dx dy=\int\limits_{0}^{\infty} \frac{1}{1+y^2} dy=tan^{-1} \infty=\frac{\pi}{2} \]

OpenStudy (anonymous):

oh very well thank u

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

@mukushla